During 2020, Martinez Company started a construction job with a contract price of $1,600,000. The job was completed in 2022. The following information is available.

2020

2021

2022

Costs incurred to date

$405,900 $830,680 $1,074,000

Estimated costs to complete

584,100 262,320 –0–

Billings to date

302,000 898,000 1,600,000

Collections to date

272,000 818,000 1,420,000

Compute the amount of gross profit to be recognized each year, assuming the percentage-of-completion method is used.

Gross profit recognized in 2020

Gross profit recognized in 2021

Gross profit recognized in 2022
Answer #1

Answers

  • Requirement asked

2020

2021

2022

Gross Profits (Loss)

$250,100

$135,220

$140,680

  • Working

Working

2020

2021

2022

A

Contract Price

$1,600,000

$1,600,000

$1,600,000

B

Cost Incurred to Date

$405,900

$830,680

$1,074,000

C

Estimated cost yet to be incurred to complete the contract

$584,100

$262,320

$0

D = B+C

Total Cost

$990,000

$1,093,000

$1,074,000

E = (B/D) x 100

% of Completion

41.00%

76.00%

100.00%

F = A x E

Revenue to date

$656,000

$1,216,000

$1,600,000

G

Revenue of Previous year

$0

$656,000

$1,216,000

H = F - G

Net Revenue this year

$656,000

$560,000

$384,000

I (=B)

Cost to date

$405,900

$830,680

$1,074,000

J

Cost to date of previous year

$0

$405,900

$830,680

K = I - J

Net Cost for the year

$405,900

$424,780

$243,320

L = H - K

Gross Profits

$250,100

$135,220

$140,680